A lottery has a very large number of tickets, one in every 500 of which entitles the purchaser to prize. Calculate the minimum no. of tickets the agent must sell to have 95% chance of selling at least one prize winning ticket.


1
Expert's answer
2021-03-11T13:47:12-0500

The probability that ticket entitle the purchaser to prize is "p = \frac{1}{500} = 0.002"

If purchaser buy N tickets, the probability that none of them will be winning according to the formula of probability of independent events intersection:

"p_{lose} = (1 - p)^N"

So, the probability that purchaser has at least one prize winning ticket is:

"p_{win} = 1 - p_{lose} = 1 - (1-p)^N"

By condition, this value should be not less than "p_0 = 0.95" :

"p_0 \leq 1 - (1-p)^N => N \geq log_{1-p}(1 - p_0) = log_{0.998}0.05 = 1496.36"

So, the minimum no. of tickets the agent must sell to have 95% chance of selling at least one prize winning ticket is 1497.